OpenStudy (dengeki_daisy):

A polyhedron has 6 vertices and 9 edges. How many faces does it have? 3 5 7 9

563blackghost (563blackghost):

Use Eulear's Formula... \(\Huge{F+V=E+2}\) Input and then simplify...

OpenStudy (dengeki_daisy):

what???

563blackghost (563blackghost):

\(\Large{F}\)ace \(\Large{V}\)ertices \(\Large{E}\)dges We input what we know of which is the vertices and the edges... \(\Huge{F+6=9+2}\) We need to simplify to solve for \(\Large{E}\)... \(\Huge\color{purple}{F+6=11 \rightarrow \color{red}{ 11-6=F}}\)

OpenStudy (dengeki_daisy):

OHHHH, 5!!

563blackghost (563blackghost):

Yup

OpenStudy (dengeki_daisy):

I totally get it now!!! This is sooo easy ^-^ Thank you